Show that the relation R on Z × Z defined by (a, b) R (c, d) if and only if a + d = b + c


is an equivalence relation.


Note: A relation on a set A is called an equivalence relation if it is reflexive, symmetric,


and transitive.


1
Expert's answer
2022-01-04T15:12:12-0500

A relation R on "Z \times Z" defined by "(a, b) R(c, d)" iff "a+d=b+c."

we see that,

"a+b=b+a \forall a, b \in z"

Therefore "(a, b) R(a, b) \forall(a, b) \in R"

therefore R is reflexive relation

Let (a, b) R(c, d)

"\begin{aligned}

&\Rightarrow a+d=b+c \\

&\Rightarrow c+b=d+a \\

&\Rightarrow(c, d) R(a, b)

\end{aligned}"

Therefore R is symmetric relation.

"\begin{aligned}

&\text{Let, (a, b) R(c, d) and (c, d) R(a, f)}\\

&\Rightarrow a+d=b+c and c+f=d+e\\

&\Rightarrow(a+d)+(c+f)=(b+c)+(d+e)\\

&\Rightarrow(a+f)+(d+c)=(b+e)+(d+c)\\

&\Rightarrow a+f=b+e\\

&\Rightarrow(a, b) R(e, f)\\

\end{aligned}"

Therefore R is transitive relation.

Therefore R is an equivalence relation.

since, "\quad a \rightarrow d=b+c"

"\Rightarrow d=-a+b+c"

Therefore "[(a, b)]=\{(c,-a+b+c) ; c \in \mathbb{Z}\}"